OpenStudy (anonymous):

please solve whether y is a function of x 25(x-10)² y² = 81x² .please help!!!!

OpenStudy (anonymous):

y = +/- (9x / 5(x-10)) i think it is not .. consider y^2 = x y = +/- sqrt(x) so when x = 5 for example you get two different y .. which is not a function the same goes here..

OpenStudy (anonymous):

x(y) you must put it in the form x=... (25x^2-500x-2500)y^2=81x^2 you will not be able to pull x all alone

OpenStudy (anonymous):

\[ 25(x-10)² y² = 81x² \Rightarrow y = \sqrt{\frac{(9x)^2}{(5(x-10))^2}} \Rightarrow y = \frac{(9x)}{5(x-10)}\] I think this implies \[ y= f(x)= \frac{(9x)}{5(x-10)} \] so y seems to be a function of x

OpenStudy (anonymous):

@tutordiffeq: y is a function of x implies \[ y=f(x) \] Check here: http://www.missouriwestern.edu/cas/Math/Functions.pdf

OpenStudy (anonymous):

FoolForMath@ you forgot that it might be +/-.. so it is not a function of x!

OpenStudy (anonymous):

@Coolsector: I think ... \[ \large \sqrt{x^2} = (x^2)^{\frac{1}{2}} = x\] Hence,...